A neophyte player catches a 125 g ball moving at 25.0 m/s in 0.02 s. A professional player catches the same ball in 1.0 s by slightly retracting his hand during the catch. Find the forces exerted by the ball on the hands of the two players.


1
Expert's answer
2021-01-23T15:52:43-0500

"F=\frac{ p}{\Delta t}=\frac{mv}{\Delta t},"

"F_{np}=\frac{0.125\cdot 25}{0.02}=156.25 ~N,"

"F_{pp}=\frac{0.125\cdot 25}{1}=3.13~N."